OpenStudy (anonymous):

Two functions are shown in the table below. Function 1 2 3 4 5 6 f(x) = −x2 + 4x + 12 g(x) = x + 8 Complete the table on your own paper, then select the value that is a solution to f(x) = g(x). x = 3 x = 4 x = 5 x = 6

OpenStudy (anonymous):

@phi

OpenStudy (welshfella):

just plug in the values into the 2 functions and put them into the table

OpenStudy (welshfella):

e. g. for f(x) x = 3 = -3^2 +4*3 + 12 = -9 + 12 = 15 g(x) x = 3:- g(3) = 3 + 8 = 11

OpenStudy (welshfella):

note f(3) and g(3) are a short way of saying the values of f(x) and g(x) when x = 3

OpenStudy (welshfella):

when x is a value such that f(x) = g(x) that is the solution